OpenStudy (anonymous):

Are 24, 45, and 51 the side measurements of a special right triangle? Yes or No, Explain your answer and show work.

OpenStudy (anonymous):

Yes, because 24^2 + 45^2 = 51^2 which fits the form of a^2 + b^2 = c^2 where "c" is the hypotenuse.

OpenStudy (anonymous):

Additionally, the relationship a^2 + b^2 = c^2 is characteristic of right triangles.

OpenStudy (anonymous):

So, the side with measurement "51" is the hypotenuse and the other two sides are the legs.